# Aa Article Title

### v1

@@ Appendix B Proof of Proposition 3
The first factor $\left(d_{1}d_{2}n2^{-i}\right)$ comes from the search of all samples at each coordinate combination in each node at this level. The second factor come from setting $R=n2^{-i}$ and $K=2$ in Lemma 2. Summing over all $\log k$ levels, the overall complexity is $\mathcal{O}\left(\log k\cdot n^{2}d_{1}^{2}d_{2}^{2}\cdot\left(\min(d_{1}d_{2},n)+\min(nd_{1},nd_{2},d_{2},d_{1})\right)\right)$. Then we can simplify the term $\min(nd_{1},nd_{2},d_{2},d_{1})$ to $\min(d_{2},d_{1})$ since we know that $n\geq 1$ by assumption. ∎
Appendix C Proof of Lemma 2
@@ Appendix D Proof of Theorem 5
for the chosen loss function $\mathcal{L}$ (e.g., (3) and (4)) in the parent node $t$. When we split node $t$ into left and right nodes
@@
Appendix E Algorithms for Ensembles of Trees
@@
Input: Tensor data $\bm{X}\in\mathbb{R}^{n\times p_{1}\times p_{2}\times\ldots\times p_{d}}$, Target values $y\in\mathbb{R}^{n}$, Number of trees $T$;
for $t\leftarrow 1$ to $T$ do
       $(X_{t},y_{t})\leftarrow\text{Bootstrap sample from }(X,y)$;
       $T_{t}\leftarrow\text{Train a decision tree on }(X_{t},y_{t})$ with random subsampling of predictors at root node;
       Store all observations in each leaf of $T_{t}$;
       Fit the chosen mean/CP/Tucker models at each of the leaf nodes of $T_{t}$;
end for
$\bm{y}_{i}^{pred}=\sum_{t=1}^{T}F_{t}(\bm{X}^{(n)}[i,:,:])$;
@@
Function GeneralizedBoostingRegressor($\bm{X}\in\mathbb{R}^{n\times d_{1}\times d_{2}}$, $\bm{y}\in\mathbb{R}^{n}$, $M$, $\eta$, $p_{resample}$):
begin
       Initialize $F_{0}(x)$ to the mean of $\bm{y}$;
      for $m=1$ to $M$ do
             Compute residuals: $r_{im}=y_{i}-F_{m-1}(\bm{X}^{(n)}[i,:,:])$ for all $i$;
            if $p_{resample}>0$ then
                   Sample indices according to weights: $\bm{I}_{resample}\sim\text{Sample}(\text{Indices},\text{Weights},\text{Size}=n\times p_{resample})$;
                   $\bm{h}_{m}$ = fit($\bm{X}[\bm{I}_{resample},:,:]$, $r_{im}[\bm{I}_{resample}]$);
                   Update weights: $\text{Weights}=\text{Weights}\times\exp(\left|r_{im}\right|)$;
                   Normalize weights: Weights=Weights/$=\text{Weights}/\sum\text{Weights}$ Weights;
            else
             end if
            for each sample $i$ do
                   $\bm{y}_{i}^{pred}=F_{m-1}(\bm{X}^{(n)}[i,:,:])+\eta\bm{h}_{m}(\bm{X}^{(n)}[i,:,:])$;
             end for
            Update $F_{m}(x)=F_{m-1}(x)+\eta\bm{h}_{m}(x)$;
       end for
      return $F_{M}(x)$;
end

### v2

@@ F Proof of Proposition 1
The first factor $\left(d_{1}d_{2}n2^{-i}\right)$ comes from the search of all samples at each coordinate combination in each node at this level. The second factor come from setting $R=n2^{-i}$ and $K=2$ in Lemma 2. Summing over all $\log k$ levels, the overall complexity is
$\mathcal{O}\left(\log k\cdot n^{2}d_{1}^{2}d_{2}^{2}\cdot\left(\min(d_{1}d_{2},n)+\min(nd_{1},nd_{2},d_{2},d_{1})\right)\right).$
Then we can simplify the term $\min(nd_{1},nd_{2},d_{2},d_{1})$ to $\min(d_{2},d_{1})$ since we know that $n\geq 1$ by assumption. ∎
G Proof of Lemma 2
@@ H Proof of Theorem 3
for the chosen loss function $\mathcal{L}$ (e.g., (9)) in the parent node $t$. When we split node $t$ into left and right nodes
@@
$\left\langle u,v\right\rangle_{t}\coloneqq\frac{1}{N_{t}}\cdot\sum_{\bm{X}^{(n)}[i,:,:]\in t}u(\bm{X}^{(n)}[i,:,:])\cdot v(\bm{X}^{(n)}[i,:,:])$
@@
I Algorithms for Ensembles of Trees
@@
2 Input: Tensor data $\bm{X}\in\mathbb{R}^{n\times p_{1}\times p_{2}\times\ldots\times p_{d}}$, Target values $y\in\mathbb{R}^{n}$, Number of trees $T$;
3 for $t\leftarrow 1$ to $T$ do
4       $(X_{t},y_{t})\leftarrow\text{Bootstrap sample from }(X,y)$;
5       $T_{t}\leftarrow\text{Train a decision tree on }(X_{t},y_{t})$ with random subsampling of predictors at root node;
6       Store all observations in each leaf of $T_{t}$;
7       Fit the chosen mean/CP/Tucker models at each of the leaf nodes of $T_{t}$;
8
9 end for
[…]
11
@@ 7       Fit the chosen mean/CP/Tucker models at each of the leaf nodes of $T_{t}$;
2
3Function GeneralizedBoostingRegressor($\bm{X}\in\mathbb{R}^{n\times d_{1}\times d_{2}}$, $\bm{y}\in\mathbb{R}^{n}$, $M$, $\eta$, $p_{resample}$):
4begin
5       Initialize $F_{0}(x)$ to the mean of $\bm{y}$;
6
7      for $m=1$ to $M$ do
8             Compute residuals: $r_{im}=y_{i}-F_{m-1}(\bm{X}^{(n)}[i,:,:])$ for all $i$;
9
10            if $p_{resample}>0$ then
11                   Sample indices according to weights: $\bm{I}_{resample}\sim\text{Sample}(\text{Indices},\text{Weights},\text{Size}=n\times p_{resample})$;
12                   $\bm{h}_{m}$ = fit($\bm{X}[\bm{I}_{resample},:,:]$, $r_{im}[\bm{I}_{resample}]$);
13                   Update weights: $\text{Weights}=\text{Weights}\times\exp(\left|r_{im}\right|)$;
14                   Normalize weights: Weights=Weights/$=\text{Weights}/\sum\text{Weights}$ Weights;
15
16            else
17
18             end if
19            for each sample $i$ do
20                   $\bm{y}_{i}^{pred}=F_{m-1}(\bm{X}^{(n)}[i,:,:])+\eta\bm{h}_{m}(\bm{X}^{(n)}[i,:,:])$;
21
22             end for
23            Update $F_{m}(x)=F_{m-1}(x)+\eta\bm{h}_{m}(x)$;
24
25       end for
26      return $F_{M}(x)$;
27
28 end
29
@@ 23            Update $F_{m}(x)=F_{m-1}(x)+\eta\bm{h}_{m}(x)$;
J Training time for fixed $N$, increasing $d_{2}$
Here we perform an experiment to illustrate how the training time of our and other tensor methods increases as tensor dimension $d_{2}$ increases under a fixed value of the sample size $N=100$ and tensor dimension $d_{1}=25$. The data is generated in the same way as in the experiments in Figure 5 of the main text.
Figure 8 shows that as tensor dimension $d_{2}$ increases, the GP-based methods (TensorGP and GPST) have the largest increase and total training time, followed by the TT methods, then by CP, then by Tucker. (All TT models with the same max_depth have roughly the same training time, so for each depth we show only TT_CP without pruning. Also, we show only max_depth=3,6 to reduce visual clutter.) These results complement and are similar in spirit to the findings from our analysis of Figures 5 and 6 in the main text.
Figure 8: In-sample and out-of-sample training time (seconds) against the second tensor dimension $d_{2}$ for the experiments in Figure 5. Here the sample size is $N=100$ and the first tensor dimension is $d_{1}=25$. For TT models, we experiment with max_depth=3,6 and select $\alpha=0.5$ in (17) and perform only one fit on an AMD Ryzen 5 3600XT 6-core, 12-thread processor.
K Out-of-sample MSE with entrywise input noise
Here we perform an experiment to illustrate the robustness of our and other tensor methods to independent entrywise noise in the input tensors. The experiment includes various levels of input noise; we set the standard deviation of each entry’s noise equal to the value input_noise_sd times the input tensor’s standard deviation.
Figure 9 shows that TT_CP and TT_Tucker with pruning are robust to input_noise_sd up to $0.3$, regardless of the max_depth. We also see that without pruning, TT_CP and TT_Tucker perform dramatically worse as depth increases, which indicates that the pruning procedure is correctly decreasing the depth of the tree to mitigate overfitting the input noise. Once input_noise_sd equals 1, all methods (except for TT_mean with pruning, which has a large test MSE even with no input noise) display a noticeable increase in test MSE, indicating that TT_CP and TT_Tucker with pruning are among the most robust methods tested in this experiment.
Figure 9: Test MSE for the experiments in Figure 5 for input tensors of size $(N,d_{1},d_{2})=(100,25,25)$. Zero-mean Gaussian noise is added entrywise to the input tensors so that each model is provided only noisy versions of the input tensors. The noise standard deviation is input_noise_sd multiplied by the input’s standard deviation. For TT models, we experiment with max_depth=1,3,5 and select $\alpha=0.5$ in (17) and perform 20 replicates.
L MSE comparison to existing forest methods
This section explores the performance of regression models — Fréchet Random Forest (FRF, bootstrap ratio = 0.75 and other parameters are default values), global Fréchet regression (GRF), and Tensor Tree (TT) with gradient boosting (GB) — on synthetic datasets. For code implementation, pyfrechet (https://github.com/matthieubulte/pyfrechet) by Bulté and Sørensen (2024) only supports 2-mode tensor as a Fréchet space input and scalar outputs, since their focus is spherical data. FrechForest (CRAN package, last maintained 2019, used by Capitaine et al. (2024); Qiu et al. (2024)) supports mainly curves as Fréchet space input and scalar output, but also supports 2- and 3-mode tensors. In contrast, our open-source implementation supports up to 4-mode tensor inputs and outputs, with improved efficiency, significantly extending the usability of the existing tensor nonparametric regression software.
The first experiment explores different noise levels and number of trees in the ensemble methods. The target variable is $y=2\cdot X[:,0]\cdot X[:,2]+3\cdot X[:,1]\cdot X[:,2]\cdot X[:,3]+\varepsilon$, where $\varepsilon$ is Gaussian noise with a configurable noise variance. The input $X$ is an $n=1000,d=5$ matrix with i.i.d. standard uniform entries, and we embed $X$ as a 3-tensor (adding an all-zero third dimension to be compatible with TT implementation) tailored to simulate structured data scenarios. For the parameter settings of the tested methods, pyfrechet’s FRF has no option for limiting each tree’s maximum depth, so we set the minimum number of observations in each leaf node to two. For FrechForest’s FRF we use the default parameter settings other than the number of trees to grow. The TT models use max_depth=4, split_rank=4 and full exploration rate $\alpha=1$, and we also execute TTentrywise GB ensembles (with maximum depth 4 and at least two observations per leaf node) in order to compete with the two mentioned ensemble methods. (If desired, Algorithm 5 shows how to train a random forest using TT.) The models are trained and evaluated over 10 random seeds to ensure robustness in MSE (on a held-out test data) and time measurements shown in Table 4. Comparing single-tree models, TT has competitive performance and fast fits compared to either FRF implementation with one tree. We expect that TT will outperform FRF for higher mode tensor inputs for a similarly simple input-response relationship, since higher mode tensors require more samples to have accurate estimates of Fréchet distances. Comparing ensemble methods, TT GB with 5 or 50 trees has larger test MSEs than FRF with the same number of trees. However, FRF seems to improve very little past 50 trees, whereas TT GB with 500 trees has a much smaller test MSE than FRF with any of the tested number of trees. Thus, of all tested ensemble methods, TT GB is able to provide by far the smallest test MSE for this scenario.
method # trees noise variance 0 noise variance 0.01 noise variance 0.1
TT 1 0.108 (1s) 0.089 (1s) 0.138 (1s)
TTentrywise GB $5$ 1.136 (12s) 1.135 (11s) 1.138 (9s)
TTentrywise GB $50$ 0.510 (33s) 0.521 (41s) 0.516 (42s)
TTentrywise GB $250$ 0.039 (196s) 0.038 (221s) 0.038 (204s)
TTentrywise GB $500$ 0.015 (588s) 0.014 (574s) 0.019 (557s)
GFR n/a 0.258 (1s) 0.194 (1s) 0.199 (2s)
pyfrechet FRF $1$ 0.098 (4s) 0.077 (4s) 0.114 (5s)
pyfrechet FRF $50$ 0.055 (24s) 0.045 (19s) 0.067 (15s)
pyfrechet FRF $250$ 0.052 (111s) 0.043 (109s) 0.066 (114s)
pyfrechet FRF $500$ 0.047 (246s) 0.049 (245s) 0.068 (241s)
FrechForest FRF $1$ 0.145 (2s) 0.150 (2s) 0.176 (2s)
FrechForest FRF $5$ 0.054 (4s) 0.055 (4s) 0.071 (4s)
FrechForest FRF $50$ 0.033 (23s) 0.033 (23s) 0.044 (23s)
FrechForest FRF $250$ 0.031 (108s) 0.031 (108s) 0.042 (108s)
FrechForest FRF $500$ 0.031 (210s) 0.030 (209s) 0.042 (210s)
Table 4: Out-of-sample MSEs and fit-predict time (seconds) for each forest method with different number of trees in the ensemble method.
The second experiment compares TT with one tree against (FrechForest’s) FRF for higher mode tensor inputs and a more complex input-response relationship. To be compatible with FrechForest, we reduce the three channels to one in the simulation code in the experiments in Figure 5 of the main text in order to reduce the number of modes of the input tensors from 4 to 3. This 3-tensor input has dimensions $n=750$ and $d_{1}=d_{2}=25$. The test MSE is computed on a held-out set with $n=250$. Figure 10 shows that with pruning, TT_CP and TT_Tucker with max_depth 1 have a similar test MSE to FRF with at least five trees but fit much faster (FRF with 250 trees took around 11 minutes to fit). These two TT methods also have about half of the test MSE and half of the training time than FRF with one tree. In this scenario, TT with one tree is faster to fit and has competitive performance to FRF with any number of trees.
Figure 10: Test MSE for modified experiments in Figure 5 for input tensors of size $(N,d_{1},d_{2})=(1000,25,25)$. The number after the “Depth_” label indicates the max_depth of the TT method. The number after the “FRF_” label indicates the number of trees. All models are run with 20 replicates. Each FRF with 250 trees took about 11 minutes (660 seconds) to fit.
M RPE results comparing the rrr method to TT
2 143 ± 9 136 ± 9 218 ± 3 223 ± 4 172 ± 2 3581 ± 79
3 172 ± 9 162 ± 7 219 ± 3 219 ± 3 132 ± 2 3633 ± 84
4 177 ± 10 171 ± 9 219 ± 3 219 ± 3 134 ± 2 3678 ± 90
5 178 ± 5 170 ± 5 216 ± 3 216 ± 3 133 ± 2 3628 ± 92
6 174 ± 7 207 ± 24 219 ± 4 219 ± 4 137 ± 2 3715 ± 106
7 260 ± 17 319 ± 24 219 ± 3 219 ± 3 138 ± 2 3659 ± 85
(a) Linear. $(d_{1},d_{2})=(3,4)$. $p_{1}=15$. $\bm{y}[:,i]=\begin{cases}\bm{X}[:,0,1]+\bm{X}[:,1,1]&\text{if }i\mod 3=0\\ \bm{X}[:,1,1]+\bm{X}[:,2,0]&\text{if }i\mod 3=1\\ \bm{X}[:,2,2]+\bm{X}[:,0,3]&\text{if }i\mod 3=2\end{cases}$.
2 219 ± 5 209 ± 3 205 ± 2 206 ± 3 297 ± 2 1249 ± 18
3 231 ± 8 211 ± 4 205 ± 2 206 ± 3 260 ± 2 1265 ± 25
4 240 ± 9 218 ± 6 205 ± 2 206 ± 3 231 ± 3 1290 ± 27
5 241 ± 11 222 ± 8 205 ± 2 206 ± 3 200 ± 2 1279 ± 12
6 257 ± 8 240 ± 8 203 ± 2 204 ± 2 169 ± 2 1328 ± 23
7 288 ± 31 276 ± 32 205 ± 2 206 ± 3 261 ± 34 1332 ± 38
(b) Non-linear. $(d_{1},d_{2})=(3,4)$. $p_{1}=6$. $\bm{y}[:,i]=\sin(\bm{X}[:,i\mod 3,i\mod 4])$.
2 342 ± 69 326 ± 65 455 ± 36 464 ± 35 392 ± 43 204 ± 37
3 433 ± 91 404 ± 86 456 ± 36 465 ± 34 395 ± 42 204 ± 37
4 531 ± 106 490 ± 97 456 ± 36 466 ± 34 398 ± 42 205 ± 38
5 609 ± 115 553 ± 107 456 ± 36 468 ± 34 400 ± 42 204 ± 37
6 666 ± 127 596 ± 117 456 ± 36 468 ± 34 403 ± 42 206 ± 38
7 666 ± 127 596 ± 117 456 ± 36 468 ± 34 406 ± 42 205 ± 38
(c) Exact CP (4). $(d_{1},d_{2})=(12,6)$. $p_{1}=7$. $\bm{y}[:,i]=\tilde{\bm{X}}[:,0,1]^{2}-\bm{X}[:,0,0]$($\tilde{\bm{X}}$ is rank-4 CP reconstructed $\bm{X}$).
2 376 ± 114 335 ± 98 452 ± 42 463 ± 41 395 ± 47 208 ± 57
3 480 ± 134 427 ± 122 451 ± 41 464 ± 42 395 ± 46 208 ± 57
4 537 ± 153 491 ± 141 451 ± 41 464 ± 42 397 ± 46 209 ± 57
5 609 ± 177 545 ± 153 451 ± 41 464 ± 42 400 ± 46 206 ± 56
6 668 ± 189 591 ± 166 451 ± 41 464 ± 42 403 ± 46 206 ± 56
7 668 ± 189 591 ± 166 451 ± 41 464 ± 42 406 ± 46 209 ± 57
(d) Exact Tucker (4, 4, 4). $(d_{1},d_{2})=(12,6)$. $p_{1}=7$. $\bm{y}[:,i]=\tilde{\bm{X}}[:,0,1]^{2}-\bm{X}[:,0,0]$ ($\tilde{\bm{X}}$ is rank-(4,4,4) or rank-(4,4) Tucker reconstructed $\bm{X}$).
Table 5: Out-of-sample RPE (multiplied by 1000 to highlight decimal differences) of various tensor-output regression models trained on synthetic data. Bold font indicates the smallest RPE of a given rank and signal; we use $\pm$ to represent the standard error over 10 repeats.
N Regression coefficients on an EEG dataset
Figure 11: Tensor-input decision tree regression coefficients on an EEG dataset in Zeng et al. (2021). First row: TRES.TRR(1D/FG/OLS/PLS) tensor-response models (following the example in $\mathtt{TRES}$) Second row: TRES.TPR(1D/FG/OLS/PLS) tensor-input models. Third row: coefficients from CP ($R=5$), Tucker ($R=5$), Fourth row: coefficients from a tensor-input tree model with depth 3 and leaf node CP models $m_{j}$ ($R=5$). Fifth row: coefficients from a tensor-input tree model with depth 3 and leaf node Tucker models $m_{j}$ ($R=5$).
[…]